OpenStudy (anonymous):

write an equation in the form y =mx+b for the line that passes through (1,-10) and has a slope of 3

OpenStudy (anonymous):

ok so m= slope so first plug in 3 for m

OpenStudy (anonymous):

whats your new formula

OpenStudy (anonymous):

what do i have to subtract to get my formula

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

just substitute 3 for m which is the slope

OpenStudy (anonymous):

(3,y1

OpenStudy (anonymous):

having hard time need to know more

OpenStudy (anonymous):

y =mx+b is your formula just put 3 where m is to get y=3x+b get it?

OpenStudy (anonymous):

ok get it

OpenStudy (anonymous):

now, take your point (x,y) which is (1,-10) and plug it in that formula that i just gave you

OpenStudy (anonymous):

10=3(1)+b

OpenStudy (anonymous):

solve for b

OpenStudy (anonymous):

3x1=3

OpenStudy (anonymous):

not quite, subtract 3 from ten, remember? to solve a variable isolate it by putting it on one side of equation with everything else on the other

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

yes thats b, no use m and b in the original formula y=mx+b

OpenStudy (anonymous):

*now, not no sorry

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

it said im wrong

OpenStudy (anonymous):

what did you put in

OpenStudy (anonymous):

y=3x+-7

OpenStudy (anonymous):

sorry, its y=3x+7 10-3=7 i didn't notice the neg sign

OpenStudy (anonymous):

i guess subtract 3 from ten confused you.... that means take 3 away from ten not ten from 3

OpenStudy (anonymous):

so sorry

OpenStudy (anonymous):

its ok you help i appreate it